Suppose $1 = 120 yen in New York, $1 = 2 euros in London, and one euro = 75 yen in Tokyo. Describe how investors use arbitrage to take advantage of the difference in exchange rates.
Answer #1

Start with 1$

Convert to euro in London=1$*eur/$ exchange rate=1*2=2 euro

Convert to yen in Tokyo=2 eur*yen/eur exchange rate=2*75=150 yen

Convert to dollar in New York=yen/(yen/$ exchange rate)=150/120=1.25 $

Started with 1$ and ended with 1.25 $..Profit of 0 25$ per 1$